# Why Agentic Theorem Prover Works: A Statistical Provability Theory of Mathematical Reasoning Models

### v1

@@ 8 Variable number of goals: truncation and tail control (compactness-preserving treatment)
The measure encoding naturally supports variable goal counts via $\mu(K)=|\text{goals}|$. However, compactness requires a mass bound.
@@ 8.1 Truncated state space and overflow event
Let $W>0$ and define $\mathcal{X}:=M_{\leq W}(K)$. Consider the overflow event
$\mathsf{Overflow}:=\left\{\max_{t\leq B}\mu_{t}(K)>W\right\}.$
@@
$\Pr^{\pi}_{\mu_{0}}(\mathsf{Overflow})\leq\delta_{W},$
@@
Couple the true and truncated processes until the first overflow time. They agree on the complement of $\mathsf{Overflow}$. On $\mathsf{Overflow}$, truncated success is $\leq$ true success (since truncation forces failure). Hence the difference is at most $\Pr(\mathsf{Overflow})$. ∎
Consequence. All bounds proved on the compact space $M_{\leq W}(K)$ extend to the original system with an additive $\delta_{W}$ (for lower bounds) or $\delta_{W}$ (for upper bounds), depending on the direction.
9 Representation learning via certificate gap minimization
@@ 9.2 The “best certificate” within a restricted function family
$U_{B,\phi}^{*}(\mu_{0})$
$:=\inf\left\{U_{B}(\mu_{0}):(U_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ U_{b+1}\geq TU_{b},\ U_{0}\geq\mathbf{1}_{G}\right\},$
$L_{B,\phi}^{*}(\mu_{0})$
$:=\sup\left\{L_{B}(\mu_{0}):(L_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ L_{b+1}\leq TL_{b},\ L_{0}\leq\mathbf{1}_{G}\right\}.$
@@ 9.3 Certificate-gap minimization objective
10 Discussion: Easy vs hard instances through the lens of provability bounds
The bounds in Sections˜6 and 7 show that the achievable success probability is governed by a small set of parameters: (1) effective horizon / proof length ($B$ or an effective $L_{\mathrm{eff}}$), (2) approximation error $\epsilon_{\mathrm{app}}$, (3) statistical error $\varepsilon_{\mathrm{stat}}(n)$ driven by geometry (doubling dimension, Lipschitz constants), (4) margin parameters $(C_{\Delta},\beta)$ for top-kk/beam stability, (5) overflow tail $\delta_{W}$.
@@ 10 Discussion: Easy vs hard instances through the lens of provability bounds
$\mathbb{E}_{\mu_{0}\sim Q}[V_{B}^{\pi_{h,k}}(\mu_{0})]$
$\quad\gtrsim\ \mathbb{E}_{\mu_{0}\sim Q}[V_{B}^{*}(\mu_{0})]-L_{\mathrm{eff}}\cdot(\epsilon_{\mathrm{app}}+\varepsilon_{\mathrm{stat}})^{\beta+1}$
$\phantom{\quad\gtrsim\ }-\delta_{W}-\text{(beam/top-$k$ bias terms)}.$
@@ 10.1 Easier instances (high success probability)
Low geometric complexity of reachable states: small $d_{\mathcal{D}}$ (local low-dimensional structure), improving $\varepsilon_{\mathrm{stat}}$.
@@ 10.2 Harder instances (low success probability)
High geometric complexity: large $d_{\mathcal{D}}$ or near-discrete geometry (huge covering numbers), inflating $\varepsilon_{\mathrm{stat}}$.
Out-of-distribution (OOD): the relevant domain $\mathcal{D}$ is poorly covered by training data, inflating both $\epsilon_{\mathrm{app}}$ and $\varepsilon_{\mathrm{stat}}$.
@@ Acknowledgments
This work was supported by JST BOOST JPMJBY24E2, JST CREST JPMJCR25I5, and JSPS KAKENHI 24K21316,
@@ Appendix C Proof of Theorem˜2
  • Base: $V_{0}^{*}(\mu)=\mathbf{1}[\mu\in G]$ is upper semicontinuous because $G$ is closed.
  • Induction step: assume $V_{b-1}^{*}$ is upper semicontinuous and bounded. For any sequence $(\mu_{n},a_{n})\to(\mu,a)$, Assumption˜1 implies $P(\cdot\mid\mu_{n},a_{n})\Rightarrow P(\cdot\mid\mu,a)$. By the Portmanteau theorem (applied to the bounded upper semicontinuous function $V_{b-1}^{*}$),
    $\limsup_{n\to\infty}\int V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu_{n},a_{n})\leq\int V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a),$
@@
  • Finally, measurable selection provides a Borel measurable maximizer $\mu\mapsto a_{b}^{*}(\mu)$.
@@ D.4 If the margin condition fails: what can we still say?
Beam/top-$k$ can still help when margins fail, but then the analysis must track set coverage rather than “gap-based stability.” A common alternative is to bound success in terms of the probability that at least one of the retained actions is “good enough,” which leads to different regularity assumptions (e.g., mass of $\epsilon$-optimal actions) rather than gap tails.
@@ F.1 Setting
    $\epsilon_{\mathrm{app}}:=\inf_{h\in\mathcal{H}}\ \|h-Q\|_{\infty}.$
@@ F.2 Sampling assumption (coverage / generative access)
independent across $j$ given $z_{i}$. (In practice, $Y_{i,j}$ can be a rollout success indicator from $(\mu,a)=z_{i}$.)
@@ F.3 Constructing an estimator
Define $\hat{h}$ as a Lipschitz extension of the values $\{\hat{q}_{i}\}$ onto $\mathcal{D}$ with Lipschitz constant at most $L_{H}$ (e.g., a McShane extension). Then $\hat{h}\in\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$. If we further project $\hat{h}$ onto $\mathcal{H}$ (or assume $\mathcal{H}$ contains such extensions), we can ensure $\hat{h}\in\mathcal{H}$ up to $\epsilon_{\mathrm{app}}$.
@@ F.4 Statement (Uniform sup norm bound; explicit form)
$\|\hat{h}-Q\|_{\infty}\leq\epsilon_{\mathrm{app}}+(L_{H}+L_{Q})\eta+\sqrt{\frac{\log(2N/\delta)}{2m}}.$ (2)
@@
$\|\hat{h}-Q\|_{\infty}\ \lesssim\epsilon_{\mathrm{app}}+(L_{H}+L_{Q})^{\frac{d}{d+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d+2}}.$
@@
Taking supremum over $z\in\mathcal{D}$ and using Equation˜3 yields the bound for the extension estimator. If $\hat{h}$ is then projected into $\mathcal{H}$, an additional $\epsilon_{\mathrm{app}}$ term appears by triangle inequality, giving Equation˜2. ∎

### v2

@@ 8 Variable number of goals: truncation and tail control (compactness-preserving treatment)
The measure encoding naturally supports variable goal counts via $x(\mathcal{G})=|\text{goals}|$. However, compactness requires a mass bound.
@@ 8.1 Truncated state space and overflow event
Let $W>0$ and define $\mathcal{X}:=M_{\leq W}(\mathcal{G})$. Consider the overflow event
$\mathsf{Overflow}:=\left\{\max_{t\leq B}X_{t}(\mathcal{G})>W\right\}.$
@@
$\mathbb{P}^{\pi}_{x_{0}}(\mathsf{Overflow})\leq\delta_{W},$
@@
Couple the true and truncated processes until the first overflow time. They agree on the complement of $\mathsf{Overflow}$. On $\mathsf{Overflow}$, truncated success is $\leq$ true success (since truncation forces failure). Hence the difference is at most $\mathbb{P}(\mathsf{Overflow})$. ∎
Consequence. All bounds proved on the compact space $M_{\leq W}(\mathcal{G})$ extend to the original system with an additive $\delta_{W}$ (for lower bounds) or $\delta_{W}$ (for upper bounds), depending on the direction.
9 Discussion on representation learning via certificate gap minimization
@@ 9.2 The “best certificate” within a restricted function family
$U_{B,\phi}^{*}(\mu_{0})$ $:=\inf\left\{U_{B}(\mu_{0}):(U_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ U_{b+1}\geq TU_{b},\ U_{0}\geq\mathbf{1}_{G}\right\},$
$L_{B,\phi}^{*}(\mu_{0})$ $:=\sup\left\{L_{B}(\mu_{0}):(L_{b})_{b=0}^{B}\subset\mathcal{F}_{\phi},\ L_{b+1}\leq TL_{b},\ L_{0}\leq\mathbf{1}_{G}\right\}.$
@@ 9.3 Certificate-gap minimization objective
10 Discussions on scaling laws
Even without cut, the bounds already yield clean scaling relationships.
10.1 Estimation-error scaling on doubling domains
From Theorem˜6, with doubling dimension $d_{\mathcal{D}}$,
$\varepsilon_{\mathrm{est}}(n)\ \approx\ n^{-1/(d_{\mathcal{D}}+2)}\cdot\mathrm{polylog}(n).$
10.2 Regret scaling (worst-case vs margin)
Worst case (Theorem˜4): if $\varepsilon_{b}\approx\varepsilon$,
$V_{B}^{*}(x_{0})-V_{B}^{\pi_{h}}(x_{0})\ \lesssim\ B\varepsilon.$
Under a margin condition (Theorem˜5):
$\mathbb{E}[V_{B}^{*}-V_{B}^{\pi_{h,k}}]\ \lesssim\ B\varepsilon^{\beta+1}.$
Plugging $\varepsilon\approx\varepsilon_{\mathrm{app}}+\varepsilon_{\mathrm{est}}(n)$ yields sample-complexity scaling for achieving a target regret $\eta$:
If $\varepsilon_{\mathrm{app}}$ is negligible and we use worst-case regret:
$Bn^{-1/(d_{\mathcal{D}}+2)}\lesssim\eta\quad\Rightarrow\quad n\gtrsim(B/\eta)^{d_{\mathcal{D}}+2}.$
Under margin:
$Bn^{-(\beta+1)/(d_{\mathcal{D}}+2)}\lesssim\eta\quad\Rightarrow\quad n\gtrsim(B/\eta)^{(d_{\mathcal{D}}+2)/(\beta+1)}.$
10.3 Where “exponential improvements” can come from
The above formulas show that performance depends critically on: the effective horizon $B$ (or a smaller $L_{\mathrm{eff}}$), the effective dimension $d_{\mathcal{D}}$, and the margin exponent $\beta$.
Therefore, any structural mechanism (e.g., cut-introduction) that: (1) reduces $B$ to $L_{\mathrm{eff}}\approx O(\log\mathrm{Complexity})$, or (2) reduces $d_{\mathcal{D}}$ by localizing to a low-dimensional manifold, or (3) increases $\beta$ by making correct actions more separated, can change required sample sizes from $\mathrm{poly}(\mathrm{Complexity})$ to $\mathrm{polylog}(\mathrm{Complexity})$, which is a standard sense in which one observes exponential-scale improvements.
11 Discussions on easy vs hard instances
The bounds in Sections˜6 and 7 show that the achievable success probability is governed by a small set of parameters: (1) effective horizon / proof length ($B$ or an effective $L_{\mathrm{eff}}$), (2) approximation error $\varepsilon_{\mathrm{app}}$, (3) statistical error $\varepsilon_{\mathrm{est}}(n)$ driven by geometry (doubling dimension, Lipschitz constants), (4) margin parameters $(C_{\Delta},\beta)$ for top-kk/beam stability, (5) overflow tail $\delta_{W}$.
@@ 11 Discussions on easy vs hard instances
$\mathbb{E}_{x_{0}\sim Q}[V_{B}^{\pi_{h,k}}(x_{0})]\ \gtrsim\ \mathbb{E}_{x_{0}\sim Q}[V_{B}^{*}(x_{0})]-L_{\mathrm{eff}}\cdot(\varepsilon_{\mathrm{app}}+\varepsilon_{\mathrm{est}})^{\beta+1}-\delta_{W}-\text{(beam/top-$k$ bias terms)}.$
@@ 11.1 Easier instances (high success probability)
Low geometric complexity of reachable states: small $d_{\mathcal{D}}$ (local low-dimensional structure), improving $\varepsilon_{\mathrm{est}}$.
@@ 11.2 Harder instances (low success probability)
High geometric complexity: large $d_{\mathcal{D}}$ or near-discrete geometry (huge covering numbers), inflating $\varepsilon_{\mathrm{est}}$.
Out-of-distribution (OOD): the relevant domain $\mathcal{D}$ is poorly covered by training data, inflating both $\varepsilon_{\mathrm{app}}$ and $\varepsilon_{\mathrm{est}}$.
@@ Acknowledgments
This work was supported by JST BOOST JPMJBY24E2, JST CREST JPMJCR25I5, and JSPS KAKENHI 24K21316
@@ Appendix C Proof of Theorem˜2
  • Base: $V_{0}^{*}(x)=\mathbf{1}[x\in G]$ is upper semicontinuous because $G$ is closed.
  • Induction step: assume $V_{b-1}^{*}$ is upper semicontinuous and bounded. For any sequence $(x_{n},a_{n})\to(x,a)$, Assumption˜1 implies $P(\cdot\mid x_{n},a_{n})\Rightarrow P(\cdot\mid x,a)$. By the Portmanteau theorem (applied to the bounded upper semicontinuous function $V_{b-1}^{*}$),
    $\limsup_{n\to\infty}\int V_{b-1}^{*}(x^{\prime})P(dx^{\prime}\mid x_{n},a_{n})\leq\int V_{b-1}^{*}(x^{\prime})P(dx^{\prime}\mid x,a),$
@@
  • Finally, measurable selection provides a Borel measurable maximizer $x\mapsto a_{b}^{*}(x)$.
@@ D.4 If the margin condition fails: what can we still say?
Beam/top-$k$ can still help when margins fail, but then the analysis must track set coverage rather than “gap-based stability.” A common alternative is to bound success in terms of the probability that at least one of the retained actions is “good enough,” which leads to different regularity assumptions (e.g., mass of $\varepsilon$-optimal actions) rather than gap tails.
@@ F.1 Setting
    $\varepsilon_{\mathrm{app}}:=\inf_{h\in\mathcal{H}}\ \|h-Q\|_{\infty}.$
@@ F.2 Sampling assumption (coverage / generative access)
independent across $j$ given $z_{i}$. (In practice, $Y_{i,j}$ can be a rollout success indicator from $(x,a)=z_{i}$.)
@@ F.3 Constructing an estimator
Define $\hat{h}$ as a Lipschitz extension of the values $\{\hat{q}_{i}\}$ onto $\mathcal{D}$ with Lipschitz constant at most $L_{H}$ (e.g., a McShane extension). Then $\hat{h}\in\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$. If we further project $\hat{h}$ onto $\mathcal{H}$ (or assume $\mathcal{H}$ contains such extensions), we can ensure $\hat{h}\in\mathcal{H}$ up to $\varepsilon_{\mathrm{app}}$.
@@ F.4 Statement (Uniform sup norm bound; explicit form)
$\|\hat{h}-Q\|_{\infty}\leq\varepsilon_{\mathrm{app}}+(L_{H}+L_{Q})\eta+\sqrt{\frac{\log(2N/\delta)}{2m}}.$ (2)
@@
$\|\hat{h}-Q\|_{\infty}\ \lesssim\varepsilon_{\mathrm{app}}+(L_{H}+L_{Q})^{\frac{d}{d+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d+2}}.$
@@
Taking supremum over $z\in\mathcal{D}$ and using Equation˜3 yields the bound for the extension estimator. If $\hat{h}$ is then projected into $\mathcal{H}$, an additional $\varepsilon_{\mathrm{app}}$ term appears by triangle inequality, giving Equation˜2. ∎